# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v3

@@ 1. Introduction
Figure 1. The proportion of private transactions
in DeFi attacks between 2022 and 2023.
@@
Key Observations. Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services. Moreover, as shown in Figure 2, the proportion of adversarial transactions sent using private services exhibits a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023. We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers. Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor function during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims. As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$. Based on the aforementioned studies and observations, we design our method for preventing DeFi attacks via unveiling adversarial contracts.
@@
To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our attack contract dataset. And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration. We extract a selection of features from the contracts capturing patterns exhibited during both their implementation and their deployment stages. Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance. The selection methods and criteria for building this dataset are described in details in section 4. These data have been made publicly available to facilitate future research in this domain.
@@
  • A custom designed smart contract bytecode lifting and analysis pipeline with integrated semantic tokenization protocol that is able to efficiently extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
@@
After providing some background in Section 2, we present an overview of LookAhead in Section 3. Section 4 details our dataset building strategies and Section 5 explains our methods, Section 6 reports our experimental results. After a brief discussion in Section 7 and 8, we conclude the paper in Section 9.
@@ 2.2. Maximal Extractable Value (MEV)
Private Mempool Services. In response to MEV, researchers and industry practitioners have proposed various countermeasures (Yang et al., 2022). The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022). This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others in the public mempool. However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
@@ 3.1.2. Adversarial Patterns
While DeFi attack logic can vary based on in the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives. Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls. Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows. Based on these observations, ML techniques can be employed to effectively identify adversarial contracts — regardless of the specific attack type — by learning and predicting these common patterns.
@@ 3.2.2. Adversaries
In a prior research (Parhizkari et al., 2023), it was observed that 79.49% of attacks in their dataset exposed the victim contract address before the adversarial transaction. Using the same method, we found 202 attacks (53.9%) with similar exposure behaviour in our dataset of $375$ adversarial contracts. Hence, we make one additional assumption about $\mathcal{A}$:
@@ 3.2.4. Rescue Time Frame
Our threat model leaves a short rescue time frame for detection and prevention to take place by the defenders and the targets, and enables the possibility of on-time reaction by victims of ongoing attacks. The rescue window is defined by $t^{window}=t_{first}-t_{deploy}-t^{pred}$, it has a typical value observed in our dataset of $t^{window}\geq 100$. Within the rescue time frame several counterattack actions could be taken, we describe those in more details in §3.3.
@@ 3.3. System Model
Defense Mechanisms. After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners. 1) If the victims’ addresses were found in the adversarial contract, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts. Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner. 2) We plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses. If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
@@ 5.2.1. Implementation Feature Extraction
To enable the extraction of implementation features (cf. §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse. First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$. Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers) in combination with the PublicFunctionSelector relation, which maps public functions to their selectors, to determine the total number of public and private functions. This combination also helps identify flashloan callback functions. Finally, to analyze the external function calls within a contract, we design a rule called ExternalCallResolved. This rule provides details about all resolved external calls made within the contract, including their selectors and signatures. Additionally, we implemented logic to handle special cases, such as identifying fallback functions and transfers of native tokens, based on the arguments and values passed to the calls. With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
@@ 5.2.2. CFG Construction and Pruning
Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements. Functions have 4-byte signatures derived from the Keccak hash (SHA3) of their canonical names and can be either private or public. Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements. Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9. The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B=\{\mathcal{BB}^{i}_{j}:\mathcal{BB}^{i}_{j}\in\mathcal{F}_{i}\}$ where $\mathcal{F}_{i}\in F$, and $(\mathcal{BB}^{i}_{x},\mathcal{BB}^{i}_{y})$ is an arc in $A_{i}$ if $\mathcal{BB}^{i}_{x}$ is a predecessor of $\mathcal{BB}^{i}_{y}$ and $\mathcal{BB}^{i}_{y}$ is a successor of $\mathcal{BB}^{i}_{x}$. Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names. Gigahorse can only recover the names for some of the public functions and external call target methods, hence if the derivation of canonical names from function signatures failed for some public function $\mathcal{F}^{public}_{i}\in F^{public}$, then $\mathcal{F}^{public}_{i}$ will be given a name of $UnknownFunction_{i}$. For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$. Next, we remove all statements not related to function calls from every block $\mathcal{BB}^{i}_{j}\in B_{i}$. Then, we perform a depth-first search (DFS) over $G$ to remove any block $\mathcal{BB}^{i}_{k}$ that contains no statements (i.e. $|\mathcal{BB}^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours. Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
@@ 5.2.3. Semantic Recovery
Similar to our fund source labelling algorithm, we attempt to match the target address with an address associated with a known label in Etherscan’s address label cloud and public name tags. For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label. Contract developers might choose to store the target address in the EVM storage or as a variable/constant which affects our ability to directly perform address matching. We use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage. If this process did not result in a matching label, we assign the address a default label UnknownTarget.
Function signatures are generated by hashing the function names with a secure hashing algorithm, and the hashing process is a one-way trap door. However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings. Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their Application Programming Interfaces (API) upon failed recovery by Gigahorse.
@@ 5.3.2. Ensemble Classifier
We use K-Nearest Neighbors (KNN) as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space. KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
@@ 6.1.1. Dataset
A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment. Contracts deployed over a different periods of time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve and change over time as well. One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks. Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation. Instead, to achieve such an objective, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing. Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
@@ 6.2.2. Experiment Results
Table 1 presents our experiment results for evaluating the performance of LookAhead’s ML models. We calculate the performance metrics of various models used in the process of forming the LookAhead detection system, including $4$ candidate classifiers, a transformer-based classifier, and a meta classifier. Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$). Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across all metrics compared to other candidates. Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples. Other models also perform well. Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method. The meta classifier, K-nearest Neighbors Neural Network (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
@@ 6.2.3. Comparison with the Baseline
Universality. We compare the performance of LookAhead in detecting all types of DeFi attacks with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts, and show that our system cannot not only match their efficacy in identifying reentrancy attacks, but can also offer similar effectiveness for detecting other types of attacks.
To evaluate LookAhead’s performance against reentrancy attack contracts, we tested our trained meta classifier model on all the reentrancy related adversarial contracts in our test dataset. We observe that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024).
Performance. The only similar research that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024). However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts. We use Forta as the baseline for comparing the performance of our ML models.

### v4

@@ 1. Introduction
Figure 1. Percentage of DeFi attacks using
private transactions between 2022 and 2023.
@@
Key Observations. Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services. Moreover, as shown in Figure 2, the proportion of adversarial transactions sent using private services exhibits a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023. We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers. Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims. As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$. Based on the aforementioned studies and observations, we design our method for preventing DeFi attacks via unveiling adversarial contracts.
@@
To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our attack contract dataset. And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration. We extract a selection of features from the contracts capturing patterns exhibited during both their implementation and their deployment stages. Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance.
@@
  • A smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
@@ 2.1. Ethereum & Decentralized Finance (DeFi)
Tokens. Tokens are cryptocurrencies created using smart contracts. They come in two main types: fungible tokens and non-fungible tokens (NFT). Furthermore, Ether, used as the reward for anyone contributing in the consensus, is referred to as the native token.
Decentralized Exchange (DEX). Unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange for trading. Instead, it conducts asset transfers directly on the blockchain via smart contracts, enabling transparency, security, and decentralization.
@@ 2.2. Maximal Extractable Value (MEV)
Private Mempool Services. In response to MEV, researchers and industry practitioners have proposed various countermeasures (Yang et al., 2022). The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022). This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others. However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
@@ 3.1.2. Adversarial Patterns
While DeFi attack logic can vary based on the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives. Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls. Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows. Based on these observations, ML techniques can be employed to effectively identify adversarial contracts — regardless of the specific attack type — by learning and predicting these common patterns.
@@ 3.2.2. Adversaries
In a prior research (Parhizkari et al., 2023), it was observed that 79.49% of attacks in their dataset exposed the victim contract address before the adversarial transaction. Using the same method, we found 202 attacks (53.9%) with similar exposure behaviour in our dataset of $375$ adversarial contracts. Hence, we also consider a weaker adversary $\mathcal{A}^{\prime}$ satisfying one more assumption than $\mathcal{A}$:
@@ 3.2.4. Rescue Time Frame
Our threat model leaves a short rescue time frame for detection and prevention to take place by the defenders and the targets, and enables the possibility of on-time reaction by victims of ongoing attacks. The rescue window is defined by $t^{window}=t_{first}-t_{deploy}-t^{pred}$, it has a typical value observed in our dataset of $t^{window}\leq 100$. Within the rescue time frame several counterattack actions could be taken, we describe those in more details in §3.3.
@@ 3.3. System Model
Defense Mechanisms. After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners. 1) To defend against $\mathcal{A}^{\prime}$, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts. Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner. 2) To counter $\mathcal{A}$, we plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses. If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
@@ 5.2.1. Implementation Feature Extraction
To enable the extraction of implementation features (cf. §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse. First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$. Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions. This combination also helps identify flashloan callback functions. Finally, to analyze the external function calls within a contract, we design a rule called ExternalCallResolved. This rule provides details about all resolved external calls made within the contract, including their selectors and signatures. Additionally, we implemented logic to handle special cases, such as identifying fallback functions and transfers of native tokens, based on the arguments and values passed to the calls. With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
@@ 5.2.2. CFG Construction and Pruning
Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements. Functions have 4-byte selector derived from the Keccak hash (SHA3) of their canonical names and can be either private or public. Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements. Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9. The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B=\{\mathcal{BB}^{i}_{j}:\mathcal{BB}^{i}_{j}\in\mathcal{F}_{i}\}$ where $\mathcal{F}_{i}\in F$, and $(\mathcal{BB}^{i}_{x},\mathcal{BB}^{i}_{y})$ is an arc in $A_{i}$ if $\mathcal{BB}^{i}_{x}$ is a predecessor of $\mathcal{BB}^{i}_{y}$ and $\mathcal{BB}^{i}_{y}$ is a successor of $\mathcal{BB}^{i}_{x}$. Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names. Gigahorse can only recover the names for some of the public functions and external call methods, hence if the derivation of canonical names from function selectors failed, we will leave them as-is and recover their names in the semantic recovery stage. For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$. Next, we remove all statements not related to function calls from every block $\mathcal{BB}^{i}_{j}\in B_{i}$. Then, we perform a depth-first search (DFS) over $G$ to remove any block $\mathcal{BB}^{i}_{k}$ that contains no statements (i.e. $|\mathcal{BB}^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours. Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
@@ 5.2.3. Semantic Recovery
Similar to our fund source labelling algorithm, we attempt to match the target address with an address associated with a known label in Etherscan’s address label cloud and public name tags. For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label. Contract developers might choose to store the target address in the EVM storage or as a variable/constant, in which case we use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage. If this process did not result in a matching label, we assign the address a default label UnknownTarget.
Function selectors are generated by hashing the function signatures with a secure hashing algorithm, and the hashing process is a one-way trap door. However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings. Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their APIs upon failed recovery by Gigahorse. If the recovery still fails, we fallback to the default label UnknownFunc.
@@ 5.3.2. Ensemble Classifier
We use K-Nearest Neighbors (KNN) to as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space. KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
@@ 6.1.1. Dataset
A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment. Contracts deployed over a different periods of time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve and change over time as well. One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks. Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation. Instead, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing. Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
@@ 6.2.2. Experiment Results
Table 1 presents our experiment results for evaluating the performance of LookAhead’s ML models. We calculate the performance metrics of various models used in the process of forming the LookAhead detection system, including $4$ candidate classifiers, a transformer-based classifier, and a meta classifier. Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$). Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across most metrics compared to other candidates. Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples. Other models also perform well. Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method. The meta classifier, K-nearest Neighbors (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
@@ 6.2.3. Comparison with the Baseline
Universality. We compare the performance of LookAhead with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts to demonstrate the universality of the LookAhead system. When our trained meta classifier model was tested on all the reentrancy related adversarial contracts in our test dataset, we observed that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024). This observation along with the experiment results presented in §6.2.2 show that despite being trained on contracts used in a variety of attacks of different types (cf. §4.1.1), LookAhead still achieves comparable performance in detecting any specific types of DeFi attacks.
Performance. The only similar work that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024). However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts. We use Forta as the baseline for comparing the performance of our ML models.